amyprez756:

Tanya’s swim club increased their $29 membership fee by $6 a year. What is the approximate percent increase in the membership fee?

jhonyy9:

use the rule of 3 simple $29 mean 100% $6 mean x x = 6*100/29 = ? x will be the right answer hope helped understandably

Tranquility:

While that method might not be intuitive, it does lead you to the right answer. Here is another way you can solve your question To calculate the percent change, use the following formula % change = (new value - original value)/(original value) * 100

jhonyy9:

@Tranquility do you think % change = [(35-29)/29 ]*100 = so this is exactly what i ve wrote above 6*100 /29 or not ?

Tranquility:

It leads to the same thing if you simplify it! % change = [(35-29)/29 ]*100 = 6*100 / 29